Theory and Applications of Categories, Vol. 27, 2012, page 1

INTRODUCTION

The International Category Theory Conference (CT2011) was held on the University of British Columbia campus in Vancouver, Canada, during July 17-23, 2011. There were 87 registered participants giving 54 talks, including 6 invited talks, 7 plenary talks by members of the program committee and 41 contributed talks. The Pacific Institute of Mathematics (PIMS) was a sponsor and gave partial support to the invited speakers and provided for the coffee breaks as well as printing a Conference Manual. Continuing help throughout the meetings was supplied by Oriana Bella, the PIMS coordinator.

The choice for the Wednesday afternoon excursions was between a Lynn Canyon hike led by Laura Scull and a Vancouver City Tour. These excursions were followed by a harbour dinner cruise attended by a total of 95 people, including participants and accompanying persons. The conference banquet was held at Sage Bistro on Friday the 22nd of July, 2011.

Richard Wood agreed to be the chair of the Program Committee when it was first formed in 2010 and he selected all its members. Because of Richard’s other commitments arising later that year, Stephen Lack kindly agreed to take over as chair. The other members of the committee were Michael Barr, George Janelidze, Tom Leinster, Susan Niefield, Diana Rodelo, Giuseppe Rosolini and Myles Tierney. Together they selected all the speakers and took care of program changes during the meetings.

Members of the Organizing Committee were John MacDonald (Chair), Maret Christiansen, Jorge Picado and Richard Wood.

Guest editors for this special CT2011 volume are Marino Gran, George Janelidze, Stephen Lack, John MacDonald and Walter Tholen. Special thanks go to Robert Rosebrugh for helpful comments and suggestions.

This item may be accessed at http://www.tac.mta.ca/tac/ or by anonymous ftp at

ftp://ftp.tac.mta.ca/pub/tac/html/volumes/27/intro/27-intro.{dvi,ps,pdf}

Published on 2012-03-20.

\begin{matrix}c \\ ◯ \end{matrix} Editors of Theory and Applications of Categories, 2012. Permission to copy for private use granted.

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors The typesetting language of the journal is TEX, and L^{A}TEX2e strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors

Clemens Berger, Université de Nice-Sophia Antipolis, cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com

Valeria de Paiva: valeria.depaiva@gmail.com

Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne : kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Glasgow, Tom.Leinster@glasgow.ac.uk Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Edinburgh: Alex.Simpson@ed.ac.uk James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca